# V-Max: A Reinforcement Learning Framework for Autonomous Driving

### v2

@@ 6 Conclusion
In this work, we introduced V-Max, a framework designed to make Reinforcement Learning (RL) practical for mid-to-end Autonomous Driving (AD). Built on Waymax, V-Max extends its capabilities with a JAX-based RL training pipeline, multi-dataset accelerated simulation, and comprehensive evaluation tools. Using these tools, we trained high-performing SAC agents, showing how V-Max can help advance RL research for AD. To further support progress in this area, we ensure full reproducibility by publishing our framework and benchmarks.
@@
Additionally, further research on adversarial scenario generation, could enable deeper robustness assessment of driving policies. ReGentS is a good starting point, diffuser-based methods could be an alternative approach Pronovost et al. (2023). Similarly, the development of more realistic simulation agents, as explored in the Waymo Open Sim Agents Challenge (WOSAC, Montali et al. (2023)) could improve realism of closed-loop simulators, reducing the reliance on non-reactive evaluation.
Acknowledgments
Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper. Only add this information once your submission is accepted and deanonymized. The acknowledgments do not count towards the 8–12 page limit.
@@ Acknowledgments
Appendix A Metrics
@@ A.1 V-Max Metrics
Appendix B Observation functions
In this appendix, we present and details the observation functions used in the experiments.
@@ Appendix B Observation functions
(b) Base
(c) Segment
(d) Lane
(e) Road
@@
Table 9: Observation configurations for the Base function.
General Configuration
Parameter Value Description
obs_past_steps 5 Past steps in obs
Object Features
features waypoints, vel, yaw, size, valid Object features
num_closest_obj 8 Closest objects count
Roadgraph Features
features waypoints, dir, types, valid Roadgraph features
interval 2 Waypoint sampling
max_meters 50 Max feature distance
roadgraph_top_k 200 Top K elements
meters_box front: 50, back: 5, left: 20, right: 20 Bounding box (m)
Traffic Light Features
features waypoints, state, valid Traffic light features
num_closest_tl 5 Closest traffic lights
Path Target Features
features waypoints Target path features
num_points 10 Target points count
points_gap 5 Gap between points
Table 10: Observation configurations for the Road function.
General Configuration
Parameter Value Description
obs_past_steps 5 Past steps in obs
Object Features
features waypoints, velocity, yaw, size, valid Object features
num_closest_obj 8 Closest objects count
Roadgraph Features
features waypoints, direction, valid Roadgraph features
interval 2 Waypoint sampling
max_meters 70 Max feature distance
roadgraph_top_k 200 Top K elements
meters_box front: 70, back: 5, left: 20, right: 20 Bounding box (m)
Traffic Light Features
features waypoints, state, valid Traffic light features
num_closest_tl 5 Closest traffic lights
Path Target Features
features waypoints Target path features
num_points 10 Target points count
points_gap 5 Gap between points
Table 11: Observation configurations for the Lane function.
General Configuration
Parameter Value Description
obs_past_steps 5 Past steps in obs
Object Features
features waypoints, velocity, yaw, size, valid Object features
num_closest_obj 8 Closest objects count
Roadgraph Features
features waypoints, direction, valid Roadgraph features
interval 2 Waypoint sampling
max_meters 70 Max feature distance
roadgraph_top_k 300 Top K elements
meters_box front: 70, back: 5, left: 20, right: 20 Bounding box (m)
Traffic Light Features
features waypoints, state, valid Traffic light features
num_closest_tl 5 Closest traffic lights
Path Target Features
features waypoints Target path features
num_points 10 Target points count
points_gap 5 Gap between points
Table 12: Observation configurations for the Segment function.
General Configuration
Parameter Value Description
obs_past_steps 5 Past steps in obs
Object Features
features waypoints, velocity, yaw, size, valid Object features
num_closest_obj 8 Closest objects count
Roadgraph Features
features waypoints, direction, types, valid Roadgraph features
max_meters 50 Max feature distance
meters_box front: 50, back: 5, left: 20, right: 20 Bounding box (m)
max_num_lanes 10 Maximum lanes
max_points_per_lane 20 Max points per lane
Traffic Light Features
features waypoints, state, valid Traffic light features
Path Target Features
features waypoints Target path features
num_points 10 Target points count
points_gap 5 Gap between points
Appendix C Training hyperparameters
In this appendix, we details hyperparameters used for all learning algorithms in the experiments.
@@ Appendix C Training hyperparameters
Behavioral Cloning (BC)
Parameter Value Description
Total Steps 200M Environment steps during training
Learning Rate 1e-4 Step size for optimization
Batch Size 64 Samples per gradient update
Updates/step 32 Gradient updates per step
Loss function Log_prob Cross entropy loss
SAC
Total Steps 25M Environment steps during training
Num Envs 16 Num of vectorized environments
Unroll Length 1 Unroll length
Learning Rate 1e-4 Step size for optimization
Batch Size 64 Samples per gradient update
Discount 0.99 Discount for future rewards
Entropy $\alpha$ 0.2 Entropy factor for exploration
Updates/step 8 Gradient updates per step
Buffer size 1_000_000 Size of replay buffer
Learning start 50_000 Random actions to prefill buffer
BC SAC
Total Steps 25M Environment steps during training
Num Envs 16 Num of vectorized environments
Unroll Length RL 1 Unroll length for RL rollout
Unroll Length IL 1 Unroll length for IL rollout
IL frequency 8 Frequency of imitation loss
RL Lr 1e-4 RL learning rate
IL Lr 5e-5 Imitation learning rate
Batch Size 64 Samples per gradient update
Discount 0.99 Discount for future rewards
Entropy $\alpha$ 0.2 Entropy factor for exploration
Updates/step 8 Gradient updates per step
Buffer size 1_000_000 Size of replay buffer
Learning start 50_000 Random actions to prefill buffer
PPO
Total Steps 200M Environment steps during training
Num Envs 16 Num of vectorized environments
Unroll Length 40 Unroll length
Learning Rate 1e-4 Step size for optimization
Batch Size 512 Samples per gradient update
Minibatches 16 Sub-samples per update
Discount 0.99 Discount for future rewards
Entropy $\alpha$ 0.2 Entropy factor for exploration
Updates/step 4 Gradient updates per step
GAE factor 0.95 GAE factor for loss calculation
Clip $\epsilon$ 0.2 Clipping factor

### v3

@@ 6 Conclusion
In this work, we introduced V-Max, a framework designed to facilitate Reinforcement Learning research for mid-to-end Autonomous Driving (AD). Built on Waymax, V-Max extends its capabilities with a JAX-based RL training pipeline, multi-dataset accelerated simulation, and comprehensive evaluation tools. Using these tools, we trained high-performing SAC agents, showing how V-Max can help advance RL research for AD. To further support progress in this area, we ensure full reproducibility by publishing our framework and benchmarks.
@@
Additionally, further research on adversarial scenario generation, could enable deeper robustness assessment of driving policies. ReGentS is a good starting point, diffuser-based methods is an alternative approach Pronovost et al. (2023). Similarly, the development of more realistic simulation agents, as explored in the Waymo Open Sim Agents Challenge (WOSAC, Montali et al. (2023)) could improve realism of closed-loop simulators, reducing the reliance on non-reactive evaluation.
@@
Appendix A Perfomances
We run a performance comparison between Waymax and V-Max, measuring steps per second (SPS) when replaying expert trajectories. The benchmark task involves converting expert steering and acceleration commands through Waymax’s InvertibleBicycle model while incrementally adding V-Max’s observation extraction and metrics computation within the environment step. All measurements were conducted on a single Nvidia L4 GPU with 16 CPUs and 64GB RAM. For fair comparison, Waymax’s default observation function was overridden to return a dummy array, as the full state observation typically reduces performance by 30-40%.
Table 8: SPS comparaison between Waymax and V-Max features.
Waymax V-Max
Single scenario, 64 objects
Env 623.08 615.63
Env + Obs N/A 543.55
Env + Metrics N/A 367.46
Env + Metrics + Obs N/A 342.72
Batch of 32 scenarios, 64 objects
Env 10109.73 9942.12
Env + Obs N/A 8210.96
Env + Metrics N/A 5387.38
Env + Metrics + Obs N/A 4609.01
Appendix B Metrics
@@
Appendix C Reimplementing PDM in V-Max
We first reimplement IDM Treiber et al. (2000), using the path target as a reference path, and considering the first object or red light along this path as the leading agent. This baseline is already quite strong. Then PDM-Closed Dauner et al. (2023), generates multiple trajectories by using PDM with the target path and two offset paths (±$\pm 1.0$ meter) and five desired speeds, namely $\{20\%,40\%,60\%,80\%,100\%\}$ of the speed limit.
External objects are forecasted using a constant-speed and constant-heading assumption, and each candidate trajectory is scored based on this forecast. The first action of the highest-scoring trajectory is then executed, and the process repeats at each timestep, which might explain the low comfort metric observed in Table 5.
Following Dauner et al. (2023), we introduce an emergency braking maneuver when the estimated time-to-collision (TTC) falls below 2.0s. While this improves PDM’s evaluation scores, it also makes the behavior less realistic.
Appendix D Observation functions
We present and details the observation functions used in the experiments. We tested different observations represented in Table 10. For object features, the module extracts waypoints, velocity, yaw, size, and validity information from the $n$ closest objects to the ego vehicle. Roadgraph features include waypoints, direction, speed limits, and validity data, sampled at 2-meter intervals within a maximum distance of 50 meters. The module selects the top 200 roadgraph elements within a bounding box extending 50 meters forward, 5 meters back, and 20 meters to each side. Traffic light features capture waypoints, state, and validity information for the 5 closest traffic lights. Finally, path target features consist of waypoints for the target trajectory, sampled from the SDC paths.
@@ Appendix D Observation functions
10x5 Waypoints
(b) road edge
(c) road edge, road line
(d) road edge, lane center
1x30 Goal Point
(e) road edge, goal point
(f) road edge, road line
(g) road edge, lane center
@@
Table 10: Observation configuration hyperparameter sweep results on WOMD Valid dataset. All experiments use SAC with LQ encoder and navigation reward (progression=0.2, off-route=-0.2).
Observation Config Road Safety Metrics Performance
Steps Obj Path Gap Features Collision Offroad Red-light Progress Accuracy V-Max
Pts $\downarrow$ $\downarrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\uparrow$
1 8 1 30 road_edge 1.54 0.62 0.27 165.65 97.11 0.81
1 8 1 30 road_edge+lane 1.79 1.25 0.43 179.84 96.03 0.82
1 8 1 30 road_edge+line 1.48 1.08 0.35 151.91 96.37 0.79
1 8 10 5 road_edge 1.54 0.63 0.21 169.1 96.97 0.87
1 8 10 5 road_edge+lane 1.46 1.31 0.15 173.9 96.54 0.84
1 8 10 5 road_edge+line 1.41 0.92 0.14 161.0 96.92 0.84
1 16 1 30 road_edge 1.56 0.73 0.20 163.91 96.80 0.79
1 16 1 30 road_edge+lane 1.08 0.71 0.19 142.1 97.53 0.87
1 16 1 30 road_edge+line 1.33 0.81 0.17 134.45 96.88 0.80
1 16 10 5 road_edge 1.28 0.64 0.29 160.5 97.33 0.87
1 16 10 5 road_edge+lane 1.60 0.87 0.15 161.9 96.55 0.87
1 16 10 5 road_edge+line 1.61 0.86 0.25 168.06 96.66 0.84
3 8 1 30 road_edge 1.38 0.61 0.15 161.93 97.30 0.81
3 8 1 30 road_edge+lane 1.09 0.82 0.23 150.5 97.28 0.84
3 8 1 30 road_edge+line 1.64 1.39 0.39 145.67 95.93 0.78
3 8 10 5 road_edge 1.84 0.62 0.13 172.5 96.51 0.86
3 8 10 5 road_edge+lane 1.92 1.33 0.26 174.5 96.05 0.85
3 8 10 5 road_edge+line 1.23 0.84 0.35 170.8 97.13 0.84
3 16 1 30 road_edge 1.00 1.06 0.17 154.22 97.10 0.81
3 16 1 30 road_edge+lane 1.49 1.26 0.20 151.0 96.46 0.85
3 16 1 30 road_edge+line 1.56 1.32 0.11 137.78 96.30 0.79
3 16 10 5 road_edge 1.14 0.69 0.12 139.2 97.09 0.87
3 16 10 5 road_edge+lane 1.03 0.94 0.10 142.2 97.18 0.87
3 16 10 5 road_edge+line 1.17 0.85 0.22 152.5 97.12 0.85
5 8 1 30 road_edge 1.70 0.68 0.28 167.38 96.82 0.80
5 8 1 30 road_edge+lane 1.52 1.33 0.19 169.8 96.38 0.83
5 8 1 30 road_edge+line 1.53 1.17 0.15 162.69 96.64 0.82
5 8 10 5 road_edge 1.23 0.92 0.21 166.4 97.04 0.87
5 8 10 5 road_edge+lane 1.59 0.87 0.15 155.25 96.78 0.81
5 8 10 5 road_edge+line 1.26 1.07 0.15 144.97 96.74 0.83
5 16 1 30 road_edge 1.35 0.62 0.25 158.57 97.26 0.82
5 16 1 30 road_edge+lane 1.45 1.22 0.24 160.5 96.56 0.84
5 16 1 30 road_edge+line 1.15 0.98 0.15 147.78 97.03 0.83
5 16 10 5 road_edge 1.09 0.73 0.17 153.6 97.24 0.88
5 16 10 5 road_edge+lane 1.55 1.20 0.18 181.9 96.49 0.84
5 16 10 5 road_edge+line 1.34 1.09 0.20 153.5 96.80 0.86
Appendix E Training hyperparameters
We details hyperparameters used for all learning algorithms in the experiments. Every algorithm are trained in one single NVIDIA L4.
@@ Appendix E Training hyperparameters
Parameter BC SAC BC+SAC PPO
Total Steps 200M 25M 25M 200M
Num Envs 16 16 16 16
Unroll Length 80 1 1(RL)/80(IL) 40
IL frequency - - 8 -
Learning Rate 1e-4 1e-4 1e-4/5e-5 1e-4
Batch Size 64 64 64 512
Num Minibatches - - - 16
Discount - 0.99 0.99 0.99
Entropy $\alpha$ 0.2 0.2 0.2
Updates/step 32 4 8 4
Buffer size - 1000000 1000000 -
Learning start - 50000 50000 -
GAE factor - - - 0.95
Clip $\epsilon$ 0.2
Loss function Log_prob - - -
Table 12: Reward shaping hyperparameter sweep results on WOMD Valid dataset. All experiments use SAC with LQ encoder, 5 observation steps, 8 objects, and road_edge features.
Reward Weights Safety Metrics Performance
Prog. Off-route Comfort Overspeed Collision Offroad Red-light Progress Accuracy V-Max
$\downarrow$ $\downarrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\uparrow$
0.1 -0.3 0 0 1.08 0.57 0.15 113.2 97.64 0.86
0.1 -0.6 0 0 1.22 0.63 0.13 111.1 97.06 0.85
0.1 -1.0 0 0 0.94 0.55 0.08 100.1 97.50 0.83
0.2 -0.2 0 0 1.18 0.67 0.15 155.4 97.45 0.87
0.3 -0.3 0 0 1.50 0.75 0.28 185.6 97.11 0.87
0.3 -0.6 0 0 1.44 0.67 0.32 184.7 96.94 0.87
0.3 -0.6 0 0 1.87 0.74 0.25 187.2 96.49 0.85
0.3 -1.0 0 0 1.94 0.98 0.23 181.3 96.16 0.84
0.5 -0.6 0 0 1.83 0.85 0.24 197.3 96.49 0.85
0.5 -0.6 0 0 2.26 0.71 0.18 190.1 95.42 0.81
0.5 -1.0 0 0 1.69 0.71 0.21 192.8 96.43 0.84
0.2 -0.2 0.1 -0.1 1.49 0.55 0.19 164.3 97.19 0.86
0.2 -0.2 0.1 -0.2 1.37 0.71 0.14 168.5 97.01 0.86
0.2 -0.2 0.2 -0.1 1.06 0.64 0.20 139.9 97.25 0.88
0.2 -0.2 0.2 -0.2 1.26 0.62 0.27 162.2 97.26 0.88
0.3 -0.2 0.1 -0.1 1.86 0.66 0.27 176.2 96.44 0.86
0.3 -0.2 0.2 -0.1 2.22 0.69 0.42 202.5 96.24 0.80
0.3 -0.2 0.2 -0.2 1.93 0.66 0.26 188.0 96.25 0.83
@@
Refer to caption Refer to caption Refer to caption
(a) Scene-centered view of the initial scenario.
Refer to caption Refer to caption Refer to caption